OpenStudy (anonymous):

what are the deminisions of a rectangle of maximum arear using 240m of fencing

sam (.sam.):

Perimeter=2x+2y Given Perimeter=240, 2x+2y=240 y=120-x ---------------------- Area, A=xy A=x(120-x) A=120x-x^2 \[\frac{dA}{dx}=120-2x\] \[For ~~maximum~~ area, ~\frac{dA}{dx}=0\] So, 120-2x=0 x=60 y=120-x y=120-60 y=60